### v1

# “Super-resolution” holographic optical tweezers array

Abstract
Aligning light spots into arbitrary shapes is a fundamental challenge in holography, leading to various applications across diverse fields in science and engineering. However, as the spot interval approaches the wavelength of light, interference effects among the spots become prominent, which complicates the generation of a distortion-free alignment. Herein, we introduce a hologram design method based on the optimisation of a nonlinear cost function using a holographic phase pattern as the optimisation parameter. We confirmed a spot interval of 0.952(1) µm in a $5\times 5$ multispot pattern on the focal plane of a high-numerical-aperture (0.75) objective by observing the near-infrared (wavelength: 820 nm) holographic output light from a spatial light modulator device, a result which overcomes the limitation of a few micrometres under similar conditions. Furthermore, the definition of the Rayleigh diffraction limit is refined by considering the separation of spots and the spot interval, thereby concluding the achievement of “super-resolution.” The proposed method is expected to advance laser fabrication, scanning laser microscopy, and cold atom physics, among other fields.
Introduction
Holography, which involves the control and utilisation of both the amplitude and phase of light, has provided numerous benefits—ranging from academic to practical applications—to the modern society for more than seventy years [1, 2]. The technological advancement of optical devices has accelerated the progress of holography; notably, the spatial light modulator (SLM) allows for the direct control of light wavefronts through the optical properties of liquid–crystal molecules [3]. Furthermore, the concept of a computer-generated hologram (CGH) facilitates the on-demand generation of desired light patterns by computing a phase pattern that reproduces the required output [4]. Holographic methods that modulate the light phase using SLMs are widely applied because of their high light utilisation efficiency [5, 6].
Generating light spot arrays is another application of holography, with uses in laser fabrication [7, 8], scanning laser microscopy [9], and cold atom physics. Holographic optical tweezers are vital tools for aligning atoms in an array of arbitrary shapes [10, 11, 12, 13, 14], where the quality and spacing of the spot arrays affect the reproducibility of the atom alignment and the strength of interactions between atoms. Generating short-spaced array patterns without distortions is a technical challenge; however, this endeavour addresses a general and fundamental issue in optics, namely, the realisation of super-resolution [15, 16, 17, 18].
A common indicator used to determine whether adjacent light spots can be visually distinguished in microscopy is Rayleigh’s criterion [19]. This criterion states that two spots cannot be distinguished if their spacing is less than the distance from the centre to the first minimum of the “Airy disk,” which is the diffraction pattern of a circular aperture under uniform illumination. However, Rayleigh’s criterion is, at most, semiquantitative and applies only to the observation of incoherent point sources. In fact, under Rayleigh’s criterion, the two Airy disks appear almost merged, which raises the question of whether each spot functions appropriately as an independent optical tweezer. This directs us to reconsider or modify the criteria according to the context of the application.
In this study, we propose a computer-generated hologram (CGH) design method that effectively generates a well-separated fine multispot light array. The proposed method is based on the optimisation of a nonlinear cost function, with the holographic phase pattern serving as the optimisation parameter. The cost function is defined as the correlation between a given target pattern and the output light pattern, which results from the light propagation from the phase distribution in the CGH. Additionally, we establish a quantitative super-resolution criterion for a multispot light array, based on the interrelationship between two conditions: spot spacing and separation of spots. The spot array is considered to achieve super-resolution if either the observed spot separation or the spacing is superior to that of two adjacent Airy disks aligned to satisfy the other conditions. By reproducing CGHs for a $5\times 5$ multispot pattern with near-infrared light ($\lambda=820$ nm), a spot spacing of $0.952(1)$ µm was confirmed for the focal plane of a high-numerical-aperture (NA$=0.75$) objective, which overcomes the limitation to a few micrometres under similar conditions.
Hologram design for fine spot arrays
Figure 1: Hologram design concept. a, Schematic of CGH design procedure. On the SLM plane, user-defined phase $\theta_{j}$ is attached to incident light amplitude $a_{j}$ to make output light amplitude of SLM to be $\psi_{j}=a_{j}e^{i\theta_{j}}$. Light propagation, expressed symbolically by an illustration of a lens, acts as a unitary transformation on $\psi_{j}$ to yield light amplitude on the image plane, $\Psi_{i}$. Comparison of output pattern, $I_{i}=|\Psi_{i}|^{2}$, and target pattern, $T_{i}$, provides a guide to update $\theta_{j}$. b, Holographic phase patterns for the 5$\times$5 square lattice pattern optimised by CGM using FOI (left) and RSS (right) cost function.
Figure 1a illustrates the concept of CGH design using a reflection-type SLM. For simplicity, the two-dimensional position on the SLM plane is denoted by $j$, which can be regarded as a vector index. The SLM applies a phase shift $\theta_{j}$ to the incident light amplitude $a_{j}$ at the $j$th position. Consequently, the output light amplitude from the SLM is expressed as $\psi_{j}=a_{j}\exp(i\theta_{j})$.
The position in the image plane can be specified by introducing another vector $i$ to approximately represent the light propagating from the SLM to the image plane using the Fourier transform matrix $U=\{u_{ij}\}$. Therefore, the light amplitude on the image plane $\Psi_{i}$ is related to $\psi_{j}$ as follows:
$\Psi_{i}=\sum_{j}u_{ij}\psi_{j},$ (1)
and the output pattern is $I_{i}=|\Psi_{i}|^{2}$. The CGH design is a problem of determining $\{\theta_{j}\}$ for the desired target pattern $T_{i}$ under fixed illumination conditions corresponding to $\{a_{j}\}$.
There are two methods for CGH design, which involve calculating the phase distribution $\{\theta_{j}\}$ to obtain the desired output patterns: self-consistent iterations and optimisation-based methods. A well-known example of the former is the iterative Fourier transform algorithm (IFTA) [6], which determines $\{\theta_{j}\}$ self-consistently by repeatedly performing light propagation calculations while imposing constraints on the light amplitudes on the SLM and image planes (Fig. 1a). In contrast, optimisation-based methods determine the optimal CGH according to the degree of correspondence between $I_{i}(\{\theta_{j}\})$ and $T_{i}$. Although optimisation-based methods require more computational resources, they offer higher computational stability and flexibility in designing CGHs.
The optimisation-based CGH design method utilising the analytical gradient of the cost function was first proposed by Harte [20]. In this approach, the set $\{\theta_{j}\}$ is determined by optimising the following cost function:
$f_{\text{RSS}}(\{\theta_{j}\})=\sum_{i}[I_{i}(\{\theta_{j}\})-T_{i}]^{2},$ (2)
which involves minimising the residual sum of squares (RSS) between $I_{i}$ and $T_{i}$. Note that $I_{i}$ and $T_{i}$ are normalised such that their sums equal unity, and $I_{i}$ is an explicit function of the phase distribution. This method provides holographic phase patterns similar to those obtained by IFTA. However, unique functionalities can be introduced into CGH design by selecting different cost functions [21]. In this study, we propose using the “fidelity of intensity” (FOI) as the cost function:
$f_{\text{FOI}}(\{\theta_{j}\})=-\sum_{i}\tilde{I_{i}}(\{\theta_{j}\})\tilde{T_{i}},$ (3)
where $\tilde{}$ denotes a quantity normalised with respect to its Euclidean norm.
Gradient-based optimisation methods including the conjugate gradient method (CGM) require computing the gradients of the cost function, which often becomes challenging for more complex functions than those considered in this study. Fortunately, $\partial f_{\text{FOI}}/\partial\theta_{j}$ can be explicitly expressed as follows:
$\frac{\partial f_{\text{FOI}}}{\partial\theta_{j}}=\frac{2}{\mathcal{N}_{I}}\sum_{i}(-\tilde{T}_{i}+f_{\text{FOI}}\tilde{I}_{i}){\rm Im}[u_{ij}^{*}\Psi_{j}\psi_{i}^{*}],$ (4)
where $\mathcal{N}_{I}(\equiv\sqrt{\sum_{i}I_{i}^{2}})$ represents the Euclidean norm of $\{I_{i}\}$. This expression enables efficient coding that completes large-scale optimisation in a reasonable time, typically 10–20 min, starting with an arbitrary random pattern. Figure 1b illustrates the examples of designed CGHs obtained using the FOI and RSS methods, respectively. Despite the scepticism regarding the practical effectiveness of the FOI-designed CGH (Fig. 1b left), we assure that it generally performs well. See Methods for details (of the hologram design).
Figure 2: Hologram reproduction experiments and results. a. Schematic of experimental setup for hologram reproduction. Output from fibre-coupled laser diode (LD) is collimated by an achromatic lens ($\text{L}_{1}$) and is projected through a neutral density filter (ND) and an aperture (AP) on the active area of a spatial light modulator device (SLM), on which phase distribution of a displayed holographic phase pattern is attached. A $4f$-telescope system comprising a pair of achromatic lenses ($\text{L}_{2}$ and $\text{L}_{3}$) is utilised to project the SLM plane onto the entrance pupil of an objective lens ($\text{OB}_{1}$). The reproduced images are captured on the focal plane of $\text{OB}_{1}$, referred to as the “image plane,” using a CMOS image sensor (IS) through another objective lens of identical specifications ($\text{OB}_{2}$) and an achromatic lens ($\text{L}_{4}$) functioning as a tube lens. A pinhole (P) is placed at the focus of the $4f$ system to eliminate the zeroth-order and stray components. The numerical and observed images reproduced from FOI- and RSS-designed CGHs are presented as well. b, c: numerical results for b FOI- and c RSS-designed CGHs. d, e: experimental results of d FOI- and e RSS-designed CGHs. The spot spacing is set to 1.8 pixels in a common target pattern, corresponding to $0.952(1)$ µm on the image plane. All results were obtained without any adaptive modifications to the hologram patterns.
Results
Hologram reproduction and analyses. To investigate the properties of the proposed method, the reproduced images of RSS- and FOI-designed CGHs were compared for target patterns as $5\times 5$ square spot arrays with varying spot spacing. The reproduced images of CGHs were observed using the experimental setup illustrated in Fig. 2a. Figures 2b-2e display examples of numerical and experimental images reproduced from FOI- and RSS-designed CGHs for a target spot spacing of 1.8 pixels, which corresponds to 0.9225 µm in the experimental setup (refer to Methods).
When the spot spacing of the target pattern is less than two pixels, the spots in the numerically reproduced images begin to merge for the RSS-designed CGHs, whereas the results of FOI-designed ones maintain clear separation. However, in the experimental reproduction, even the images from FOI-designed CGHs cannot maintain spot separation for target spot spacing of less than 1.8 pixels.
Although adaptive adjustment of hologram patterns is frequently incorporated into CGH designs to enhance the image quality of holographic outputs [22, 23], the current results were obtained without any post-modification. In our experiments, a $5\times 5$ square spot array was selected as the target pattern to minimise the computational load, which increases with the number of spots, and to verify the principles of the proposed method.
The mean spot spacing is 0.952(1) µm for Fig. 2d, which cannot be attained by any other existing hologram design methods under similar optical conditions. Actually, reproduced patterns on the focal plane was observed by super-resolution optical pattern measurement [24, 25] with the consequence that the spot spacing was estimated as less than 1 µm using 852 nm trapping beam with a high-NA objective (NA=0.75 in design). Preliminary experiments on fluorescence image observation also confirmed trapping of atoms on the focal plane. These results will be reported elsewhere in the future.
Table 1: Properties of reproduced spot arrays. The spatial uniformity is assessed by the relative standard deviation (SD), which is the ratio of the SD to the mean value, in the powers of the $5\times 5$ light spots that form the spot array. The power of each light spot is estimated by applying fitting analysis to an output image (Methods). The light utilisation efficiency is calculated as the ratio of the output light power (measured by a power meter through a 1-mm-diameter pinhole placed on the image plane) to the illumination power measured on the SLM plane. The light utilisation efficiency is simply stated as the sum of $I_{i}$ over the area displayed in Fig. 2 for numerically reproduced images. The value and error of a physical quantity indicate the mean value and standard error (SE), respectively, in a series of quantities evaluated numerically or experimentally from $20$ CGHs, each designed by starting from a different initial random pattern to generate a common target.
Spot uniformity Utilisation efficiency
FOI RSS FOI RSS
Numerical $1.21(4)\times 10^{-2}$ $5.6(10)\times 10^{-2}$ 0.120(3) 0.928(4)
Experimental $9.70(15)\times 10^{-2}$ $2.7(3)\times 10^{-1}$ 0.127(2) 0.64(3)
Spot uniformity and light utilisation efficiency. Considering the scope of this study, the evaluation of the reproduced images should be based on the characteristic properties of light-spot arrays. One might assume that the FOI itself serves as a measure of the quality of the reproduced images. However, FOI primarily reflects the total correspondence between the output and target patterns. Therefore, we investigated the properties of the reproduced images in terms of the uniformity of the light spot array and light utilisation efficiency.
The spot uniformity and utilisation efficiency of the images reproduced using FOI- and RSS-designed CGHs are presented in Table 1. The results reveal that the FOI method exhibits small variations in spot power but demonstrates low light utilisation efficiency, despite employing phase-only modulation. In contrast, the RSS results display the opposite behaviour. Uniformity was suppressed in both numerical and experimental results, whereas the numerical and experimental values of light utilisation efficiency were similar. Thus, the FOI method can generate high-quality spot arrays in terms of uniformity and stability, although at the expense of efficiency.
Figure 3: Spacing and separation of adjacent light spots. a, Focal pattern (top) and profile (bottom) of Airy disk [NA = 0.75(air)]. Red line in the bottom plot shows cross-sectional profile corresponding to the focal pattern, where $r_{\text{A}}$ refers to the Airy-disk radius. b, Focal patterns and profiles of incoherent sums of two Airy disks aligned at distances of $d=r_{\text{A}}$ (left) and $2r_{\text{A}}$ (right). In the bottom plots, thick green lines display the profiles of the incoherent sums, whereas thin red lines indicate those of individual Airy disks. Definitions of peak (P) and valley (V) are illustrated as well. c, (Solid green line)  The relationship between the VP ratio and spot spacing for incoherent sums of adjacent Airy disks is depicted, where the traditional Rayleigh diffraction limit corresponds to $\text{VP}=0.66$ at $d/r_{\text{A}}=1$. Open circles represent VP ratios for numerically reproduced images from FOI- (red open circles) and RSS-designed (blue open circles) CGHs, while closed circles represent those for the corresponding observed patterns, i.e., Figs. 2c (red closed circle) and d (blue closed circle), respectively. Error bars (SEs) are determined similarly to those in Table 1 but are smaller than the size of the plot symbols, except for the experimental result of an RSS-designed CGH.
Relation between spacing and separation of spot array. FOI-designed holograms can generate high-quality spot arrays with extremely narrow spacing. To evaluate the behaviour of spot arrays within the diffraction-limit regions, the focal patterns calculated under a high-NA [= 0.75 (air)] condition are displayed in Fig. 3. Note that all numerical results in Fig. 3 were calculated using vector Debye formulas [26, 27, 28, 29, 30], as the focal patterns of coherent light are affected by the vector nature of the light amplitude under NA conditions greater than 0.70. The Airy disk pattern calculated assuming circularly polarised light is illustrated in Figure 3a.
The Rayleigh-limit conditions portrayed on the left in Fig. 3b present the adjacent light spots (Airy disks) as inadequately separated, while well-separated light spots, as displayed on the right of Fig. 3b, are desirable for maintaining a high aspect ratio in applications such as laser processing and suppression of background noise in scanning laser microscopy. Therefore, a factor other than spot spacing is required for a proper evaluation of spot-array patterns while considering their applications.
The valley-to-peak (VP) ratio of a spot array along with the inverse of the peak-to-valley ratio was introduced as a metric for spot separation. A smaller VP ratio indicates a clear separation of adjacent light spots. The VP ratio is advantageous because of its availability and robustness in derivation under unfavourable conditions compared to other metrics, such as the size of each light spot. To derive the VP ratio of the reproduced spot-array image, the peak positions were determined by fitting the reproduced images (Methods). The peak value refers to the mean of $I_{i}$ at the peak positions of all the bright spots, whereas the valley value refers to the mean of $I_{i}$ at all the midpoints between the adjacent peaks. Thus, the VP ratio was evaluated as the ratio of the valley value to the peak value.
The blue (red) open circles displayed in Fig. 3c represent the VP ratios of numerically reproduced images from hologram patterns designed by the FOI (RSS) cost method, whereas the blue (red) closed circles indicate those derived from the experimental images. In terms of the VP ratio, the proposed FOI method outperforms the conventional RSS method in the range of $(1<d/r_{\text{A}}<2)$.
Discussions
The super-resolution criterion can be defined as both definite and quantitative. In considering the relationship between physical quantities involving spot separation, the traditional Rayleigh diffraction limit is established for an incoherent superposition of two Airy disks. Thus, this approach should be logically followed. We propose to determine whether a spot array exhibits “super-resolution” by examining the relationship between the spacing $d$ and the VP ratio of two incoherent Airy disks, as illustrated by the solid green line in Fig. 3(c). This criterion quantitatively delineates a limit that cannot be surpassed using incoherent light and is beneficial for discussing the separability of multispot patterns using coherent light.
The area below the green line can be regarded as a super-resolution region, where the separation of adjacent spots is superior to the criterion for fixed spot spacing, or where the spot spacing is smaller than the criterion for a fixed VP ratio. According to the extended super-resolution criterion, the numerical VP values in Fig. 3(c) suggest the potential of FOI-designed holograms to generate super-resolution light spot arrays in the regime $d/r_{\text{A}}<1.8$. In particular, superior spot separation is anticipated for $1.71.3<d/r_{\text{A}}<$ 1.7. The RSS-designed holograms generate holograms surrounding the criterion but cannot overcome these limitations. The experimental VP ratios at $d/r_{\text{A}}=1.4$ in Fig. 3(c) exhibit moderately larger values than the numerical ones, suggesting the influence of extrinsic factors in the experiments, such as residual optical aberrations. Nevertheless, we emphasise that the results of the FOI-cost method still fall within the super-resolution regime. Refer to Supplementary Information for further details.
As described in Methods, instabilities exist in the hologram design and optical reproduction, which can be distinguished by comparing the numerical and experimental results. A point of issue in this comparison is that errors in the numerical light utilisation efficiencies were similar for both the FOI and RSS methods, as listed in Table 1. Interestingly, the accuracy of the CGM procedure was the dominant contributor to these errors, which was the only common factor between the FOI and RSS methods. The CGM procedure contributes toward achieving higher numerical stability compared to self-consistent iterations like the IFTA method.
In contrast, the numerical and experimental values of light utilisation efficiency are similar for the FOI method, implying that the FOI method is tolerant to experimental disturbances concerning light utilisation efficiency, which is a global property of light spot arrays. However, this argument does not apply to spot uniformity, which exemplifies the local properties of the spot arrays. The numerical results, including both the values and errors of spot uniformity, reveal distinct differences between the FOI and RSS methods, as outlined in Table 1. Thus, spot uniformity is affected by the selection of the cost function.
However, this difference reduced marginally in the experimental results, implying a nonnegligible contribution from experimental factors such as aberrations in optical trains. The FOI method generates super-resolution spot arrays of higher quality than the RSS method, and the structural stability of the FOI method is maintained to a certain extent even in the presence of experimental disturbances.
The subsequent question is: how does the FOI method generate super-resolution spot arrays? A well-known example is the phase shift method [31] used in lithography. Although lithography differs from holography in terms of near and far-field light propagation phenomena, this technology has attempted to overcome the diffraction limits by controlling the phase distribution of coherent light. The RSS method can be classified under this type of approach owing to its high light utilisation efficiency and globally smooth CGHs, as illustrated in Fig. 1b.
In contrast, the FOI-designed CGH shown in Fig. 1b presents a phase profile with prominent high-frequency components that cause diffusive light propagation. This leads to the characteristics of the FOI method, namely, low light utilisation efficiency. Therefore, FOI-designed CGHs may function as complex amplitude modulation phase holograms [32, 33, 34, 35] to remove the undesired light components that prevent super-resolution and cause interference, which induces instability in the holographically reproduced images. Conversely, the current results indicate a limitation of phase-only modulation approaches and the necessity of removing undesired light components to achieve super-resolution spot arrays.
Before concluding the paper, we discuss the numerical properties of the FOI cost function. In CGH design problems, an “absolute” solution satisfying $I_{i}=T_{i}$ is not possible within a set of physical solutions at all instances due to the nature of light propagation. Consequently, CGHs designed using optimisation-based methods generally rely on cost functions. Here, it is essential to focus on the behaviour of the cost function rather than its superficial form. In fact, we can readily confirmed that $\sum(\tilde{I}_{i}-\tilde{T}_{i})^{2}=2-2\sum\tilde{I}_{i}\tilde{T}_{i}$ holds, indicating the equivalence of the FOI and RSS cost functions under Euclidean normalisation. Thus, one can infer that Euclidean normalisation itself plays a crucial role in generating CGHs for super-resolution spot array patterns. Although we were unable to establish a mathematically rigorous description of the significant changes in the reproduced images due to the cost functions, it is still worthwhile to discuss this briefly.
Consider a test target pattern such that $T_{i}=c>0$ only at separated $M$ points, among $N$ lattice points in the image plane ($M\leq N$). The support for $T_{i}$ is denoted by $\alpha$ as follows: $\alpha=\{\forall i|T_{i}\neq 0\}$. The target pattern satisfies $T_{i}=1/M\ (i\in\alpha)$ for the total-sum normalisation, whereas $\tilde{T}_{i}=1/\sqrt{M}\ (i\in\alpha)$ holds for Euclidean normalisation. Therefore, the FOI cost function becomes $f_{\text{FOI}}=-(1/\sqrt{M})\sum_{i\in\alpha}\tilde{I}_{i}$. However, the RSS cost provides:
$\sum_{i}\left(I_{i}-T_{i}\right)^{2}$ $=2-{\sum_{i\neq k}}^{\prime}T_{i}T_{k}-2\sum_{i\in\alpha}I_{i}T_{i}-{\sum_{i\neq k}}^{\prime}I_{i}I_{k}$
$=\left(1+\frac{1}{M}\right)-\frac{2}{M}\sum_{i\in\alpha}I_{i}-{\sum_{i\neq k}}^{\prime}I_{i}I_{k},$ (5)
where $\sum I_{i}^{2}=1-\sum_{i\neq k}^{\prime}I_{i}I_{k}$ and a similar relation for $T_{i}$ was used for the aforementioned deformations. Note that the sum of the output values over $\alpha$ commonly works to reduce both FOI and RSS cost values, despite the difference in normalisation. However, an additional factor in $f_{\text{RSS}}$ reduces the cost value: the last term of Equation (5). This term accounts for the autocorrelation of the output pattern independent of the target, and it broadens the light spots, which is favourable for reducing $f_{\text{RSS}}$.
Conclusions
This study discovered that the quality of narrow holographic optical lattice patterns could be significantly improved by utilising the FOI-cost method for hologram optimisation, as opposed to the traditional RSS-cost method. We have developed a super-resolution criterion specifically for arrays of light spots, which is based on the spacing of the spots and the VP ratio of the incoherently superimposed Airy disks. The multispot patterns generated in the present experiments satisfied this criterion, representing a key advancement for subsequent cold atom experiments employing holographic techniques [25]. The potential applications of this method are not limited to cold atoms; it also holds promise for laser processing [7, 8] and bioimaging [9]. In future, we intend to establish a comprehensive understanding of the relationship between other hologram design methods, such as IFTA and its derivatives, and optimisation-based methods. This exploration is expected to evolve, thereby expanding the potential application areas of the proposed method.